Dear Richard dear all,

please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO were:

Arcetri: A. Puglisi, P. Stefanini, S. Esposito, C. Arcidiacono, A. Riccardi, M. Xompero, R. Briguglio
LBTO: G. Brusa, T. Shih, A. Rakich

Total # of person: 9

W unit

1) A two lens system has been added to the CCD47 optical path to increase the unit FoV from 7.2 arcsec to 13.5 arcsec.
2) The LBTO laser tracker has been used to measure the relative orientation of the AGW flange (the one that will be bolted to the telescope rotator) and the W unit mirror alignment. The alignment mirror was set parallel to the AGW flange with an accuracy of about +/- 3 arcmin.
3) The equipment to align the W unit at the telescope has been successfully tested during the test with the laser tracker.

AO control SW

1) A procedure to estimate the wobble of the telescope PSF in the WFS focal plane has been written. The procedure will be used during the W unit alignment with the telescope.

LBT672a

1) The shipping frame has been packed to be stored.
2) The shell has been moved in clean tent

Other

Meeting
: Schedule definition meeting for AO system installation at the telescope
Action items:
1) Arcetri people together with G. Brusa and J. Little to define a technical schedule for the period 4th 7th of March.
2) G. Brusa provide an alignment plan for the 4D interferometer with retroreflector.
3) A,. Riccardi to contact D. Gallieni & R. Biasi to get shipping informations for retrotreflector, hub2 and LBT672a spare parts. This informations will be forwarded to L. Miglietta to write the letter for the customs.

Next update tomorrow.

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback